### v1

# GUI‑AIMA: Aligning Intrinsic Multimodal Attention with a Context Anchor for GUI Grounding

Abstract
Graphical user interface (GUI) grounding is a key function of computer-use agents, which maps natural-language instructions to actionable screen regions. Existing approaches based on Multimodal Large Language Models (MLLMs) typically formulate it as a text-based coordinate generation task, yet directly generating precise coordinates from visual inputs remains challenging and computationally intensive. An intuitive way to implement GUI grounding is to first select visual patches relevant to the instructions and then determine the precise click location within those patches. Based on the observations that general MLLMs have some native grounding capability, nested within their attentions, we propose GUI-AIMA, an attention-based and coordinate-free supervised fine-tuning framework for efficient GUI grounding. GUI-AIMA aligns the intrinsic multimodal attention of MLLMs with patch-wise grounding signals. These signals are calculated adaptively for diverse user instructions by multi-head aggregation on simplified query-visual attention matrices. Besides, its coordinate-free manner can easily integrate a plug-and-play zoom-in stage. GUI-AIMA-3B was trained with only 85k screenshots, demonstrating exceptional data efficiency and verifying that light training can trigger the native grounding capability of MLLMs. It achieves state-of-the-art performance among 3B models, attaining an average accuracy of 58.6% on ScreenSpot-Pro and 62.2% on OSWorld-G. Project page: https://github.com/sjz5202/GUI-AIMA.
1 Introduction
Graphical User Interface (GUI) (Hong et al., 2024; Cheng et al., 2024) agents have emerged as pivotal tools in automating interactions with digital devices, spanning from mobile applications (Rawles et al., 2023; 2024; Wang et al., 2024a; Ye et al., 2025a) to desktop software (Zhang et al., 2024; Qin et al., 2025; Deng et al., 2023; Xie et al., 2024; OpenAI, 2025; Zheng et al., 2024). GUI grounding plays an important role in mapping natural language instructions to specific elements on the screen, such as buttons, text fields, or icons (Yang et al., 2024). This process ensures that the agent’s actions are precise and contextually relevant, especially in environments with high-resolution displays and intricate layouts. However, the diversity of human instructions and GUI designs across platforms and applications (Li et al., 2025) makes GUI grounding a challenging task. Conventional methods rely on structured representations, such as HTML for web pages or accessibility trees (Zhang et al., 2025) for mobile applications. While these provide detailed information about the interface elements, they come with limited accessibility and excessive verbosity, which can lead to inefficiencies in processing. Moreover, structured data may omit essential visual cues such as layout and icons, which are critical for accurate grounding.
When humans use computers, they first determine the general area and then use visual feedback to interact with GUI elements (Ye et al., 2025b). Directly outputting coordinates from MLLMs might not be the most intuitive method. Instead, models should mimic human behavior:
Figure 1: An example of GUI-AIMA with optional two-step GUI grounding for high-res screenshots.
first selecting the relevant visual patch and then deciding the specific click position within that patch, i.e., the coordinates in the GUI. Previous studies, such as TAG (Xu et al., 2024a) and GUI-Actor (Wu et al., 2025), have explored the coordinate-free GUI grounding where corordinates are determined by selecting the most relevant visual patches in a screenshot instead of directly generating text tokens. However, GUI-Actor introduces extra modules into the MLLM, which requires an additional adaptation stage. Xu et al. (2024a) considers a vanilla but inaccurate way to identify and aggregate attention maps from all query text tokens, referred to as vanilla attention grounding.
In this work, we propose GUI-AIMA, an attention-based method for coordinate-free GUI grounding. We train the MLLM’s multi-head self-attention (MHSA) (Vaswani et al., 2017) with patch-wise grounding supervision using a novel attention head weighting mechanism. To simplify the impractical aggregation over all query tokens in vanilla attention grounding, we append a learnable <ANCHOR> token after the visual and text query tokens and supervise its attention over the visual tokens as a surrogate aggregator for all query tokens’ text-visual attentions. For precise multi-head aggregation, we weight each attention head by the magnitude of its query–visual correlation estimated from the MLLM’s visual-sink query tokens, rather than including visual-irrelevant query tokens as in vanilla attention grounding. The MLLM’s general visual-sink query tokens is identified by computing similarity between query and visual token embedding from layer-wise hidden states instead of relying on correlations within each individual attention head. These visual-sink query tokens guide the attention head weighting, prioritizing attention heads with strong interactions between visual-sink query and visual tokens, consistent with the general query-visual pattern revealed in hidden states, while down-weighting other heads. Together, compared with the vanilla attention grounding, GUI-AIMA achieves 4.5% improvements on ScreenSpot-pro. By operating directly on intrinsic attention without task-specific modules, GUI-AIMA enables data-efficient visual grounding training with only 85k screenshots. In addition, the center-encouraging and overlap-aware patch-wise labeling further boosts the performance of GUI-AIMA. Furthermore, offset grounding errors are prevalent on high-resolution screenshots due to dense visual content and detail loss from image compression in MLLMs. GUI-AIMA, with strong base grounding capacity for precise candidate-region localization, can combine the initially predicted region with an extra zoomed-in step without additional training to self-correct offset errors, as shown in Fig. 1.
The main contributions of GUI-AIMA are summarized as follows:
  • We introduce GUI-AIMA, an attention-based, coordinate-free framework that aligns intrinsic MHSA with patch-wise supervision by simplifying the vanilla attention-based visual grounding with anchored-attention predictions and designing an overlap- and center-aware patch-wise labeling scheme that better aligns with human click behaviors.
  • We propose a simple and effective attention head weighting mechanism using visual-sink query tokens to emphasizes heads exhibiting strong query-visual interactions, improving efficiency and generalization without extra grounding modules.
  • With only one-stage fine-tuning on a small open-sourced training dataset, GUI-AIMA-3B outperforms all models of similar sizes and rivals much larger MLLM-based GUI grounding models. Ablations demonstrate the benefits of the special anchored token, head-wise weighting mechanism using visual-sink query tokens, and weighted visual patch labels. GUI-AIMA also provides insights into how to understand and specialize the functionality of intrinsic multimodal attention for visual grounding.
2 Related Work
Coordinate-based GUI grounding:
The core challenge for GUI agents (Wang et al., 2024b) is grounding: aligning user instructions with the correct actionable elements in the screenshot, due to the semantic inconsistency of layouts and UI elements across diverse GUI environments (Liu et al., 2025). In early attempts, along with the screenshots, extra structured inputs, such as HTML for U-Ground (Gou et al., 2024), or extractions from visual parsing modules, such as element extractions from OmniParser (Wan et al., 2024; Yu et al., 2025) and generated captions as contexts in Aria-UI (Yang et al., 2024), are fed into MLLMs as the supplementary inputs for the better interface understanding and more precise localization. Later works, such as AGUVIS (Xu et al., 2024b), SeeClick (Cheng et al., 2024) and OS-Atlas (Wu et al., 2024), explore the GUI grounding leveraging MLLMs (Bai et al., 2025; Chen et al., 2024b) with screenshot-only inputs, enabling the end-to-end localization with improved scalability across diverse GUI environments. The grounding approach in these methods is to generate coordinate-based click centers or bounding boxes as text, which is indirect alignment of visual grounding requiring additional efforts, such as scaled GUI corpora (Qin et al., 2025; Chai et al., 2024) and OCR pretraining (Hong et al., 2024) for connecting coordinates with UI elements. Recent endeavors manage to address this gap from the data-intensive manner of the coordinate-based GUI grounding methods via incorporating GUI-specific grounding reward signals into RL-training (Luo et al., 2025; Lu et al., 2025; Liu et al., 2025; Tang et al., 2025) and performing autonomous GUI exploration (Fan et al., 2025; Wu et al., 2025).
Coordinate-free GUI grounding:
Recent works have begun to replace traditional coordinate‑based grounding text with patch‑level attention map predictions, where patches belonging to the correct region receive higher weight. TAG (Xu et al., 2024a) first leveraged the multi-head self-attention between GUI instructions and visual tokens in MLLMs for tuning‑free GUI grounding, showcasing the intrinsic potential of MLLMs’ attention patterns for GUI tasks. However, the generalization of TAG for novel interfaces is restricted by the backbone and by the heuristics used for selecting text tokens and attention heads. SE-GUI (Yuan et al., 2025) retains a coordinate‑based prediction manner but employs self‑attention to iteratively filter noisy training samples during training. Aside these attention-based improvements, GUI-Actor (Wu et al., 2025) introduces an embedding‑based grounding head that derives attention between input visual patch embeddings and the final hidden state of a special <ACTOR> token. Compared with GUI-AIMA, GUI-Actor adds an additional module to connect cross-layer visual and <ACTOR> embeddings, requiring an extra warm-up training phase and resulting in much lower training efficiency.
3 Methods
The core idea of GUI-AIMA is utilizing the intrinsic query-visual pattern within attention maps of MLLMs. It performs supervised fine-tuning on the multi-head self-attention matrices across layers. Intuitively, as shown in Section 3.1, we can extract the attention vector between a text token and all visual patch tokens in each attention map of MLLM as the patch-wise attention vector, where each visual patch token’s attention weight can indicate its degree of membership in the grounding region. And the tuning-free aggregation of patch-wise attention vectors between query tokens has shown the potential for text localization (Xu et al., 2024a). Existing barriers for attention-based GUI grounding are how to aggregate patch-wise vectors for each query text tokens and each attention head. Besides, annotation gap between patch-wise and original coordinate labels further leads to inaccurate groundings. In GUI-AIMA, we propose a general attention-based GUI grounding framework that simplify the query token selection in Section 3.3 and refine the head-wise weighting mechanism for the patch-wise attention vector aggregation in Section 3.4. Furthermore, as detailed in Section 3.5, we extend GUI-AIMA with a two-step zoom-in inference that mitigates offset errors in grounding. Compared with previous embedding-based method, such as GUI-Actor (Wu et al., 2025), GUI-AIMA does not add extra grounding modules to MLLMs, eliminating the extra warm-up training phase.
Figure 2: With user query $\mathcal{Q}$, screenshot patches $\mathcal{V}$ and multi-head attentions $\{\mathbf{A}^{l,h}\}_{l\in[L],h\in[H]}$ from the MLLM, the vanilla attention grounding needs additional aggregation between all query tokens’ grounding vectors. In our proposed simplified version, a special <ANCHOR> token can learn to implicitly aggregate all query tokens. Then we aggregate grounding vectors of <ANCHOR> token across layers and heads with carefully designed weights to produce the patch-wise predictions.
3.1 Preliminary: Inherent GUI Grounding within Multi-head Self-Attention
Multi-head Self-Attention (MHSA).
For MLLMs with $L$ layers and $H$ attention heads per layer, given the output embeddings $\mathbf{H}^{l-1}$ of the preceding transformer layer l1l-1 with dimension $d$, query and key are computed as $\mathbf{Q}^{l,h}=\mathbf{H}^{l-1}\mathbf{W}^{l,h}_{Q}$ and $\mathbf{K}^{l,h}=\mathbf{H}^{l-1}\mathbf{W}^{l,h}_{K}$ for each attention head $h$, $1\leq h\leq H$ , where $\mathbf{W}^{l,h}_{Q},\mathbf{W}^{l,h}_{K}\in\mathbb{R}^{d\times d_{h}}$ are parameters with dimension $d_{h}=d/H$. Then, the attention matrix $\mathbf{A}^{l,h}$ is computed as:
$\mathbf{A}^{l,h}=\operatorname{softmax}\left(\mathbf{Q}^{l,h}\mathbf{K}^{l,h^{\top}}/\sqrt{d_{h}}+\mathbf{M}\right),\quad\mathbf{M}_{ij}=\begin{cases}-\infty,&i<j\\ 0,&i\geq j\end{cases},$
where $\mathbf{M}$ is the attention mask for the causal attention $\mathbf{A}^{l,h}$.
Inherent GUI Grounding Indications in $\mathbf{A}^{l,h}$.
Given the interface image token index sequence $\mathcal{V}=[v_{1},\ldots,v_{|\mathcal{V}|}]$ which is patch-wise for MLLMs and the user query token index sequence $\mathcal{Q}=[q_{1},\ldots,q_{|\mathcal{Q}|}]$, we format the intermediate token sequence as $\mathbf{H}^{l-1}=[\mathbf{H}^{l-1}_{\mathcal{V}},\mathbf{H}^{l-1}_{\mathcal{Q}}]$ for layer l1l-1, where $\mathbf{H}^{l-1}_{\mathcal{V}}\in\mathbb{R}^{|\mathcal{V}|\times d}$ is the embedding sequence of visual patch tokens, $\mathbf{H}^{l-1}_{\mathcal{Q}}\in\mathbb{R}^{|\mathcal{Q}|\times d}$ is the embedding sequence of the user query tokens. With the self-attention matrix $\mathbf{A}^{l,h}$ computed from $\mathbf{H}^{l-1}$, the extracted patch-wise attention vector $\mathbf{A}^{l,h}_{q_{i},\mathcal{V}}$ between each query token $\mathbf{H}^{l-1}_{q_{i}}$ and the visual sequence $\mathbf{H}^{l-1}_{\mathcal{V}}$ can reflect the patch-wise query-visual correlation. Among all patches $\mathcal{V}$, the image patch $v_{i}$ with large attention value $\mathbf{A}^{l,h}_{q_{i},v_{i}}$ is more likely to contain regions related to $q_{i}$. Intuitively, the aggregation of all patch-wise attention vectors $\{\mathbf{A}^{l,h}_{q_{i},\mathcal{V}}\}_{l\in[L],h\in[H],i\in[|\mathcal{Q}|]}$ among $|\mathcal{Q}|$ text tokens and $L\times H$ attention heads can produce the inherent patch-wise GUI grounding indications:
$\hat{\mathbf{a}}=\frac{1}{L\cdot H\cdot|\mathcal{Q}|}\sum_{l,h,i}\,w_{q_{i},(l,h)}\,\mathbf{A}^{l,h}_{q_{i},\mathcal{V}}\quad\in\mathbb{R}^{|\mathcal{V}|},$ (1)
where $w_{q_{i},(l,h)}$ is the aggregation weights for each query-visual attention vector with $w_{q_{i},(l,h)}\geq 0$ and $\sum_{l,h,i}w_{q_{i},(l,h)}=1$. We denote the aggregation in Eq. 1 as the Vanilla Attention Grounding shown as “vanilla” in Fig. 2. In the previous training-free attempt (Xu et al., 2024a) based on this manner, every selected text tokens are treated uniformly and attention heads are roughly selected. For the better adaptation of an attention-based GUI grounding method, we need a simplified and effective weighting strategy via $w_{q_{i},(l,h)}$.
3.2 Coordinate-free Patch-wise Labeling
The inherent grounding of MLLMs via the query-visual attention in Eq. 1 is patch-wise instead of the pixel manner in the original coordinate-based bounding boxes. In order to enable the patch-wise grounding supervision, we first reformat the ground truth bounding box $gt^{\mathrm{bbox}}=\left[x_{1},y_{1},x_{2},y_{2}\right]$ into patch-wise label vectors $\bm{p}\in\mathbb{R}^{|\mathcal{V}|}$, where $|\mathcal{V}|$ is the patch size of the interface $I$. Image patch $v_{i}$ is positive only if there is an overlap between $gt^{\mathrm{bbox}}$ and patch $v_{i}$. Considering the border patches which are partially overlapped with $gt^{\mathrm{bbox}}$ should be less weighted and the patches of center region should be highlighted, we weight each patch with $p_{v_{i}}$ according to the overlapping ratio of itself with $gt^{\mathrm{bbox}}$ and the distance from center of patch $v_{i}$ to the center of $gt^{\mathrm{bbox}}$ with Gaussian distribution following GUI-$\text{G}^{2}$ (Tang et al., 2025). In more details, we define the overlap between positive image patch $i$ and grounding box annotation $gt^{\mathrm{bbox}}$ using intersection over union ratio, denoted as $\textbf{IoU}(v_{i},gt^{\mathrm{bbox}})$. We take its center point $\bm{\mu}_{v_{i}}=(\text{center}^{v_{i}}_{x},\text{center}^{v_{i}}_{y})$ and the center point $\bm{\mu}_{gt}=(\text{center}^{gt}_{x},\text{center}^{gt}_{y})$ of $gt^{\mathrm{bbox}}$ for the following weighting scheme:
$p_{v_{i}}=\textbf{IoU}(v_{i},gt^{\mathrm{bbox}})\cdot\mathcal{N}\left(\bm{\mu}_{v_{i}};\bm{\mu}_{gt},\bm{\Sigma}_{gt}\right),$ (2)
where $\bm{\Sigma}_{gt}=\mathrm{diag}\!\left((\sigma_{x}^{gt})^{2},(\sigma_{y}^{gt})^{2}\right)$ with adaptive 2D standard deviations as $\sigma_{x}^{gt}=\alpha\cdot\left(x_{2}-x_{1}\right)$ and $\sigma_{y}^{gt}=\alpha\cdot\left(y_{2}-y_{1}\right)$. Empirically, we set $\alpha=0.8$. In this manner, coordinate-free weighted patch label $\bm{p}=\mathrm{normalize}\big(\{p_{v_{i}}\}_{i=1}^{|\mathcal{V}|}\big)$ is overlapping-aware and encouraging center-clicking. With the patch-wise prediction $\hat{\mathbf{a}}$ in Eq. 1 and ground truth label $\bm{p}$ derived from annotations, the attention grounding loss $\mathcal{L}_{\text{Attn}}$ is defined as the KL-Divergence between $\bm{p}$ and $\hat{\mathbf{a}}$:
$)\mathcal{L}_{\text{Attn}}=\mathbb{D}_{\mathrm{KL}}(\bm{p}\,\|\,\operatorname{normalize}(\hat{\mathbf{a}}$)) (3)
3.3 Efficient Grounding with a Visual Anchor Token
The supervision on $\hat{\mathbf{a}}$ in Eq. 1 requires a carefully calibrated balance across $\mathbf{A}^{l,h}_{q_{i},\mathcal{V}}$ of each query token, where the weight of each query token $w_{q_{i},(l,h)}$ is difficult to determine. Moreover, direct supervision on the attention of query text tokens will impair the MLLM’s general capabilities (e.g., image captioning) and thus harm generalization (Nguyen et al., 2023).
To tackle this issue, we consider adding a special <ANCHOR> token into the vocabulary and place it after GUI inputs to format $\mathcal{V},\mathcal{Q},\texttt{<ANCHOR>}]$. This design simplifies the token-wise aggregation on $\{\mathbf{A}^{l,h}_{q_{i},\mathcal{V}}\}_{l\in[L],h\in[H],i\in[|\mathcal{Q}|]}$ and disentangle general understanding functionality from GUI grounding. <ANCHOR> token serves as a bridge between user query tokens and correct grounding patches. Intuitively, for each attention head, the patch-wise attention vector $\mathbf{A}_{\texttt{a},\mathcal{V}}^{l,h}\in\mathbb{R}^{|\mathcal{V}|}$ between <ANCHOR> token and all visual tokens $\mathcal{V}$ learns an implicit aggregation of $\{\mathbf{A}^{l,h}_{q_{i},\mathcal{V}}\}_{q_{i}\in\mathcal{Q}}$ over all query tokens. With the anchored attentions, the previous aggregation via the weight $w_{q_{i},(l,h)}$ in Eq. 1 can be simplified as the multi-head aggregation:
$\hat{\mathbf{a}}=\frac{1}{L\cdot H}\sum_{l,h}\,w_{l,h}\,\mathbf{A}^{l,h}_{\texttt{a},\mathcal{V}}\quad\in\mathbb{R}^{|\mathcal{V}|},$ (4)
with the query-ignored weight vector of attention heads $\bm{w}=[\,w_{l,h}:l\in[L],\;h\in[H]\,]\in\mathbb{R}^{1\times LH}$. This simplified attention-based grounding is shown as “simplified” in Fig. 2.
Figure 3: Details about how to compute the final patch-wise prediction based on the grounding vectors of <ANCHOR> token: GUI-AIMA first specifies visual-sink query tokens via computing hidden state similarities between query tokens and visual patches; Then it computes weights $\bm{w}$ of each attention head based on visual-sink query tokens in Eq. 8; Finally, GUI-AIMA aggregates the grounding vectors of <ANCHOR> token across layers and heads in Eq. 4.
3.4 Attention Head Weighting using Visual-sink Query Tokens
In Section 3.3, we introduce a special anchor token to avoid aggregating patch-wise attention vectors over all query text tokens. In this section, we continue to discuss how to determine $w_{l,h}$, the weights of different attention heads across different layers for the multi-head aggregation. Previous works (Li et al., 2023; Clark et al., 2019) has shown the functional diversity between attention heads across transformer blocks in LLMs. In GUI grounding, we want to identify attention heads, which shows active query-visual interactions. The attention head with large attention value between visual tokens $\mathcal{V}$ and text query tokens $\mathcal{Q}$ is more relevant, while others are not. As GUI grounding in the attention manner aims to capture the query-visual correspondence between the user instruction and interface, the above strategy focusing on query-visual interactions allows the attention supervision skew towards the attention heads with large query-visual correlations and affect the neutral attention heads less, thus improving the inherent attention grounding of MLLMs without impairing the pretrained capacity, as the adaptation complies with the original attention functionalities.
A straightforward measurement of the query-visual correlations of $\mathbf{A}^{l,h}$ is the cumulative sum of all query-visual attention entries in the vanilla attention grounding:
$w_{l,h}=\sum_{q_{i},v_{i}\in\mathcal{Q},\mathcal{V}}\mathbf{A}^{l,h}_{q_{i},v_{i}}$ (5)
This strategy is under-specified as not all text tokens in the query are necessary for query-visual connections. Since <ANCHOR> is introduced in GUI-AIMA to represent the context of the whole query sequence $\mathcal{Q}$, a similar simplification towards Eq. 5 can be applied as:
$w_{l,h}=\sum_{v_{i}\in\mathcal{V}}\,\mathbf{A}^{l,h}_{\texttt{a},v_{i}},$ (6)
which is the cumulative sum of the attention entry only between visual tokens $\mathcal{V}$ and <ANCHOR> tokens. However, <ANCHOR> is not granted to function as the representative token for context summarization at the initial stage of training, as the embedding of new introduced <ANCHOR> token is random initialized. Only relying on premature <ANCHOR> token as weights in Eq. 6 will bring in noisy and biased attention aggregations. Thus, we consider adaptively selecting query token for weighting instead of using all query tokens $\{q_{i}\}$ of the given instruction or <ANCHOR> in Eq. 5 and Eq. 6. We denote the chosen token for computing $w_{l,h}$ as visual-sink query tokens $\mathcal{Q}_{s}$, which are global active tokens for connecting visual inputs and query tokens.
Visual-sink Query Tokens $\mathcal{Q}_{s}$ Selection.
For a user query sequence $\mathcal{Q}$, each text token $q_{i}$ reflects different visual correlations with visual tokens in $(l,h)$-indexed head as $\sum_{v_{i}\in\mathcal{V}}\,\mathbf{A}^{l,h}_{q_{i},v_{i}}$, the cumulative sum of $q_{i}$’s attention weights over all visual tokens. And the text tokens with the massive visual correlation inside the model vary in different user queries. Here, in order to identify the query tokens with strong visual affinity globally for the MLLM, we directly measure the Cosine similarity with intermediate hidden states $\mathbf{H}^{l}$ shown in the step 1 of Fig. 3:
$c_{q_{i}}^{l}=\sum_{v_{j}\in\mathcal{V}}\texttt{Sim}(\mathbf{H}^{l}_{q_{i}},\mathbf{H}^{l}_{v_{j}}).$ (7)
Tokens with larger $c_{q_{i}}^{l}$ has strong visual correlations in layer $l$. The reason of measuring the query-visual affinity of $q_{i}$ in layer $l$ using hidden states $\mathbf{H}^{l}$ instead of attention, i.e. $c_{q_{i}}^{l}=\sum_{h\in[H]}\sum_{v_{i}\in\mathcal{V}}\,\mathbf{A}^{l,h}_{q_{i},v_{i}}$ is that the query-visual pattern discovered in hidden states $\mathbf{H}^{l}$ is not necessary statistically prevailing among each head’s self-attention matrix, as only a smaller subset of attention heads are “semantic heads” that key on semantic functionality and representation similarity reflected in $\mathbf{H}^{l}$ (Elhelo & Geva, 2024; Olsson et al., 2022; Voita et al., 2019) (an analysis is available in Appendix A). In comparison, the measurement in Eq. 7 is straightforward to reflect visual correlations of each text token in a global and general view. For grounding, we prioritize the “semantic” minority heads showing the similar pattern discovered from hidden states $\mathbf{H}$. Here, we denote the query tokens with $\operatorname{topK}$ large $c_{q_{i}}^{l}$ as the visual-sink query tokens $\mathcal{Q}_{s}^{l}$. The general query-visual correlation pattern of the MLLM captured by $\mathcal{Q}_{s}^{l}$ derived from hidden states should be consistent in the attention head heavily weighted for grounding for grounding, i.e., $\mathcal{Q}_{s}^{l}$ should also exhibit strong visual correlation in those heads. We measure the visual correlation on $\mathcal{Q}_{s}^{l}$ tokens in the individual attention head as the cumulative sum of attention values between $\mathcal{Q}_{s}^{l}$ and $\mathcal{V}$ in $\mathbf{A}^{l,h}$, as shown in the step 2 of Fig. 3, with further normalization:
$w_{l,h}=\sum_{q^{s}_{i}\in\mathcal{Q}_{s}^{l},v_{j}\in\mathcal{V}}\,\mathbf{A}^{l,h}_{q^{s}_{i},v_{j}},\quad w_{l,h}=\frac{\exp\left(w_{l,h}\right)}{\sum_{l^{\prime}=1}^{L}\sum_{h^{\prime}=1}^{H}\exp\!\big(w_{l^{\prime},h^{\prime}}\big)}.$ (8)
That is, for semantic attention head with the strong query-visual affinity measured as Eq. 8 based on $\mathcal{Q}_{s}$, it should be emphasized using its large $w_{l,h}$ for aggregations in Eq. 4. In this way, the highlighted attention heads share the aligned MLLM’s general patterns found in the hidden state $\mathbf{H}^{l}$ as Eq. 7. In practice, we found that unifying the same $\mathcal{Q}_{s}$ for all layers’ $w_{l,h}$, that is $\forall\,l\in\{1,\dots,L\}$:
$\textbf{Global uniform: }\mathcal{Q}_{s}^{l}=\mathcal{Q}_{s}\coloneqq\operatorname*{arg\,topK}_{q_{i}\in\mathcal{Q}}(c_{q_{i}}),$ (9)
where $c_{q_{i}}=\sum_{l=1}^{L}c_{q_{i}}^{\,l}$, leads to better training stability and performance, instead of layer-wise $\mathcal{Q}_{s}^{l}$ different in each layer:
$\textbf{Layer-wise: }\mathcal{Q}_{s}^{l}\coloneqq\operatorname*{arg\,topK}_{q_{i}\in\mathcal{Q}}(c_{q_{i}}^{\,l}).$ (10)
After obtaining $\mathcal{Q}_{s}$, we compute head-wise weight $w_{l,h}$ via Eq. 8 and perform multi-head aggregation for final prediction $\hat{\mathbf{a}}$ via Eq. 4.
Soft pattern matching considering $\mathcal{Q}_{s}$ and other tokens in weighting.
Above matching between the MLLM’s general and head-wise query-visual patterns only considering representative visual-sink query tokens $\mathcal{Q}_{s}$. Alternatively, we can define the two token-wise query-visual distributions: MLLM’s general $\mathcal{D}_{g}=\operatorname{normalize}([c_{q_{i}}]_{i=1}^{|\mathcal{Q}|})$ from $c_{q_{i}}=\sum_{1}^{L}c_{q_{i}}^{l}$, and head-wise $\mathcal{D}^{l,h}_{head}=\operatorname{normalize}([d_{q_{i}}^{l,h}]_{i=1}^{|\mathcal{Q}|})$ from $d_{q_{i}}^{l,h}=\sum_{v_{j}\in\mathcal{V}}\mathbf{A}^{l,h}_{q_{i},v_{j}}$. We can then measure the closeness between $\mathcal{D}_{g}$ and $\mathcal{D}^{l,h}_{head}$ via negative Kullback-Leibler divergence as the weight $w_{l,h}$ to encourage attention head showing similar query-visual pattern with the MLLM’s general query-visual pattern:
$w_{l,h}=-\mathbb{D}_{\mathrm{KL}}(\mathcal{D}_{g}\|\mathcal{D}_{head}^{l,h}).$ (11)
This strategy provides the soft matching between all query tokens. In addition to $\mathcal{Q}_{s}$, the visual-irrelative query tokens are also included, but with less contributions. We denote this soft matching variant as GUI-AIMA-3B (soft).
3.5 Two-step Inference with Zoom-in
Under GPU memory constraints, high-resolution GUI screenshots are usually down-sampled, yielding fewer visual patch tokens for the MLLM. The resulting information loss and reduced fine-grained spatial granularity inevitably harm grounding accuracy. GUI-AIMA provides flexible spatial granularity since its patch-wise grounding. We can easily perform a two-step inference by adding zoom-in without extra training: (1) feed the compressed high-resolution screenshot to predict the approximate location. We use the center of this predicted location to determine a specific area to focus on. (2) we re-run the inference on the newly cropped region, providing a much more accurate result. This two-step inference targets to mitigates failure cases on high-resolution screens where the model identifies the right region but the center prediction is slightly offset the ground-truth box. Detailed observations and analysis of the zoom-in strategy is provided in Section 4.3.
4 Experiments
Table 1: Performance comparison of different models across various task categories based on Text, Icon, and Average scores on ScreenSpot-Pro. “-” indicates unreported results in original papers. Methods with are Qwen-2.5-VL-based.
Model CAD Dev Creative Scientific Office OS Average Text Icon Text Icon Text Icon Text Icon Text Icon Text Icon Text Icon Avg. General GPT-4o 2.0 0.0 1.3 0.0 1.0 0.0 2.1 0.0 1.1 0.0 0.0 0.0 1.3 0.0 0.8 Claude Computer 14.5 3.7 22.0 3.9 25.9 3.4 33.9 15.8 30.1 16.3 11.0 4.5 23.4 7.1 17.1 Qwen2.5-VL-3B 9.1 7.3 22.1 1.4 26.8 2.1 38.2 7.3 33.9 15.1 10.3 1.1 23.6 3.8 16.1 Qwen2.5-VL-7B 16.8 1.6 46.8 4.1 35.9 7.7 49.3 7.3 52.5 20.8 37.4 6.7 38.9 7.1 26.8 SFT OS-Atlas-7B 12.2 4.7 33.1 1.4 28.8 2.8 37.5 7.3 33.9 5.7 27.1 4.5 28.1 4.0 18.9 UGround-V1-7B 15.8 1.2 51.9 2.8 47.5 9.7 57.6 14.5 60.5 13.2 38.3 7.9 45.2 8.1 31.1 UI-TARS-72B 18.8 12.5 62.9 17.2 57.1 15.4 64.6 20.9 63.3 26.4 42.1 15.7 50.9 17.6 38.1 Jedi-3B 27.4 9.4 61.0 13.8 53.5 8.4 54.2 18.2 64.4 32.1 38.3 9.0 49.8 13.7 36.1 Jedi-7B 38.0 14.1 42.9 11.0 50.0 11.9 72.9 25.5 75.1 47.2 33.6 16.9 52.6 18.2 39.5 UI-TARS-1.5-7B 38.6 11.0 58.4 12.4 58.1 15.4 66.7 21.9 74.6 35.9 49.5 13.5 57.5 16.9 42.0 GUI-Actor-3B - - - - - - - - - - - - - - 42.2 GUI-Actor-7B - - - - - - - - - - - - - - 44.6 RL UI-R1-E-3B 37.1 12.5 46.1 6.9 41.9 4.2 56.9 21.8 65.0 26.4 32.7 10.1 - - 33.5 InfiGUI-R1-3B 33.0 14.1 51.3 12.4 44.9 7.0 58.3 20.0 65.5 28.3 43.9 12.4 49.1 14.1 35.7 GUI-G1-3B 39.6 9.4 50.7 10.3 36.6 11.9 61.8 30.0 67.2 32.1 23.5 10.6 49.5 16.8 37.1 SE-GUI-3B 38.1 12.5 55.8 7.6 47.0 4.9 61.8 16.4 59.9 24.5 40.2 12.4 50.4 11.8 35.9 GUI-G2-3B 20.3 9.4 42.2 2.1 43.4 2.8 43.8 10.0 46.3 15.1 29.0 4.5 37.6 6.0 25.5 GUI-AIMA-3B 43.7 25.0 63.0 20.0 57.1 17.5 68.1 33.6 69.5 35.9 47.7 22.5 58.1 24.2 45.2 GUI-AIMA-3B $(soft)^{*}$ 53.3 23.4 63.0 24.1 55.1 14.7 66.7 30.9 76.8 34.0 55.1 21.4 61.6 23.5 47.1 GUI-AIMA-3B + zoom-in 66.5 34.4 75.3 33.1 64.7 23.1 75.7 38.2 85.9 62.3 65.4 49.4 72.1 36.8 58.6
Implementation Details.
We apply Qwen2.5-VL-3B-Instruct (Bai et al., 2025) as the MLLM backbone for GUI-AIMA and all ablation variants. We follow the special token setting in GUI-Actor (Wu et al., 2025) and retain the next-token prediction loss for GUI-AIMA’s grounding format. The training of GUI-AIMA-3B is conducted on 8 NVIDIA A100-80G GPUs with effective global batch size 64 and learning rate 5e-6. We set $\alpha$ as 0.8 for the adaptive deviation control in Eq. 2. Regarding the setting of visual-sink query token $\mathcal{Q}_{s}$, we select the top-1 global large tokens complying with Eq. 9. As for the training dataset, we employ the entire training set of GUIAct (Chen et al., 2024a), AndroidControl (Li et al., 2024), Wave-UI (Jeffries & Team, 2024), randomly sampled 60k samples from UGround (Gou et al., 2024) and from GTA1 training set (Yang et al., 2025), respectively, with 259k instructions and $\sim$85k images in all. For the 45k ablation training set for Section 4.2, we select the first 10k samples from GUIAct, AndroidControl, Wave-UI and first 15k samples from UGround. GUI-AIMA-3B is trained for one epoch with all parameters unfrozen, without extra modules and a warmup stage. Implementation of baselines is in Appendix B. For two-step inference introduced in Section 3.5, we set the crop size to $448$ pixels and adopt a $2\times$ zoom-in ratio for the second pass.
Evaluation datasets and metrics.
We evaluate GUI-AIMA, baselines and ablation variants on ScreenSpot-v2 (Wu et al., 2024) for general GUI visual grounding evaluation across mobile, desktop and web domains, ScreenSpot-Pro (Li et al., 2025) for testing on challenging higher-resolution screenshots from diverse and complex professional software scenarios, e.g. interfaces from different Operating Systems with large distribution gap, and OSWorld-G (Xie et al., 2025a). All benchmarks contain separated text-centered and icon-centered visual grounding tasks, where the icon set is the more abstract visual-based task less hinted by text. For grounding accuracy, we follow the standard center point-based metric (Lin et al., 2024) that the correct click-point prediction locates within the ground-truth bounding box.
Baselines.
Three categories of methods are compared: (1) General MLLMs including GPT-4o (OpenAI, 2024), Gemini-2.5-Pro (Comanici et al., 2025), Claude Computer (Hu et al., 2024), Operator (OpenAI, 2025), Qwen2.5-VL (Bai et al., 2025); (2) GUI-specific supervised fine-tuned model: OS-Atlas (Wu et al., 2024), UGround (Gou et al., 2024), UI-TARS (Qin et al., 2025), JEDI (Xie et al., 2025b) and GUI-Actor (Wu et al., 2025) ; (3) GUI-specific Reinforcement fine-tuned model: UI-R1 (Lu et al., 2025), InfiGUI-R1 (Liu et al., 2025), GUI-G1 (Zhou et al., 2025), SE-GUI (Yuan et al., 2025), GUI-G2 (Tang et al., 2025).
Table 2: Performance comparison of different models across various task categories based on Mobile, Desktop, Web and Average scores on ScreenSpot-v2. “-” indicates unreported results in original papers. Methods with are Qwen-2.5-VL-based.
Model Mobile Desktop Web Avg.
Text Icon Text Icon Text Icon
General
Operator 47.3 41.5 90.2 80.3 92.8 84.3 70.5
GPT-4o + OmniParser-v2 95.5 74.6 92.3 60.9 88.0 59.6 80.7
Qwen2.5-VL-3B 93.4 73.5 88.1 58.6 88.0 71.4 80.9
Qwen2.5-VL-7B 97.6 87.2 90.2 74.2 93.2 81.3 88.8
SFT
OS-Atlas-7B 95.2 75.8 90.7 63.6 90.6 77.3 84.1
UGround-V1-7B 95.0 83.3 95.0 77.8 92.1 77.2 87.6
UI-TARS-1.5-7B 95.9 84.8 94.9 80.7 90.6 86.2 89.7
UI-TARS-7B 96.9 89.1 95.4 85.0 93.6 85.2 91.6
Jedi-3B 96.6 81.5 96.9 78.6 88.5 83.7 88.6
Jedi-7B 96.9 87.2 95.9 87.9 94.4 84.2 91.7
GUI-Actor-3B 97.3 86.9 95.5 81.7 94.6 81.4 90.4
RL
UI-R1-E-3B 83.0 97.1 85.0 91.7 77.9 95.4 89.2
GUI-G2-3B 96.5 82.5 95.4 75.0 88.4 72.4 86.3
GUI-AIMA-3B 98.5 85.8 95.6 85.7 93.6 82.3 90.8
GUI-AIMA-3B $(soft)^{*}$ 99.2 84.8 95.6 84.9 94.3 80.2 90.5
4.1 Main Results: Grounding Performance
We compare GUI-AIMA with state-of-the-art methods on ScreenSpot-Pro, ScreenSpot-v2 and OSWorld-G and report results in Table 1, Table 2 and Table 3. While only being trained with 259k samples without filtering or careful selection, GUI-AIMA-3B outperforms all same scale 3B MLLM-based models and shows comparable or even better results than larger scaled models. Among Qwen-2.5-VL based models, marked with in both tables, GUI-AIMA-3B can better handle the visual grounding in complex software scenarios across different platforms and achieves the best performance on the most challenging ScreenSpot-Pro benchmark, especially on the abstract icon task set. Specifically, GUI-AIMA-3B is better than strong large size coordinate-based UI-TARS-1.5-7B, JEDI-7B, and also better than the embedding-based coordinate-free GUI-Actor-7B model, highlighting the superiority of directly supervising on the multi-head self-attention weights instead of modeling the query-visual attention map via hidden states. On ScreenSpot-v2, GUI-AIMA-3B achieve the comparable results with the strongest baselines, such as JEDI-7B and UI-TARS-7B, and better than the same size GUI-Actor-3B. Besides the advanced performance, another advantage is the training efficiency of GUI-AIMA with only 259k training elements, as most supervised fine-tuned baselines trained on millions of GUI elements. For the comparison with reinforcement fine-tuned baselines, while both trained with smaller training sets than SFT baselines, GUI-AIMA-3B performs better and shows better generalization. Among GUI-AIMA-3B and GUI-AIMA-3B (soft), GUI-AIMA-3B (soft) handles the high-resolution interfaces in ScreenSpot-Pro with 1.9% improvments, and GUI-AIMA-3B is better on dealing with diverse graphic environments in ScreenSpot-v2 and OSWorld-G. The two-step inference with zoom-in significantly improve the performance of GUI-AIMA on high-resolution benchmarks, ScreenSpot-pro and OSWorld-G, demonstrating the flexibility of GUI-AIMA’s attention-based patch-wise grounding for inference-time improvements.
4.2 Ablations
Section 4.2 shows the ablation results on ScreenSpot-v2 and ScreenSpot-Pro. All the ablation variants are trained on the 45k ablation training data with Qwen2.5-VL-3B-Instruct as the backbone. GUI-Actor (45k) is trained in two stages same as the original setting.
Compare Coordinate-free modeling manners.
In Section 4.2, we include 3 different coordinate-free GUI grounding manners, one embedding-based: GUI-Actor which relies on hidden embeddings for computing similarity with extra modules, and two attention-based: vanilla attention grounding as Eq. 1 and simplified attention grounding as Eq. 4. With the same importance on all the query tokens and attention heads, the vanilla attention grounding can converge faster than GUI-Actor on more complex visual grounding tasks in ScreenSpot-Pro. When simplifying the aggregation on query tokens, even with naive ”weighting uniformly”, the simplified attention grounding can have 1.50% and 4.24% improvement over GUI-Actor for ScreenSpot-v2 and ScreenSpot-Pro. These results demonstrate the effectiveness and faster convergence of attention-based visual grounding methods than embedding-based methods, and shows the advantage of compressing query contexts into <ANCHOR> token instead of manually control the grounding importance from each query token.
Table 3: Performance comparison of different models across various task categories based on Mobile, Desktop, Web and Average scores on OSWorld-G. Methods with are Qwen-2.5-VL-based.
Model
Text
Matching
Element
Recognition
Layout
Understanding
Fine-grained
Manipulation
Avg.
General
Operator 51.3 42.4 46.6 31.5 40.6
Gemini-2.5-Pro 59.8 45.5 49.0 33.6 45.2
Qwen2.5-VL-3B 41.4 28.8 34.8 13.4 27.3
Qwen2.5-VL-7B 45.6 32.7 41.9 18.1 31.4
SFT
OS-Atlas-7B 44.1 29.4 35.2 16.8 27.7
UGround-V1-7B 51.3 40.3 43.5 24.8 36.4
UI-TARS-7B 60.2 51.8 54.9 35.6 47.5
Jedi-7B 65.9 55.5 57.7 46.9 54.1
GUI-Actor-7B 65.9 62.7 66.4 38.2 56.6
UI-TARS-1.5-7B 52.6 75.4 72.4 66.7 64.2
Jedi-3B 67.4 53.0 53.8 44.3 50.9
GUI-Actor-3B 64.4 60.6 64.8 33.6 54.6
GUI-AIMA-3B 62.8 64.9 68.4 32.9 56.9
[…]
GUI-AIMA-3B+ zoom-in 69.4 67.0 74.3 45.4 62.2
Table 4: Ablation results on ScreenSpot-v2 and ScreenSpot-Pro of coordinate-free GUI grounding methods fine-tuned with a fixed 45k randomly sampled dataset. Variants in blue represent the selected settings for GUI-AIMA.
[…]
Text Icon Avg. Text Icon Avg.
[…]
GUI-Actor-3B (45k) 96.24 75.99 87.42 50.67 12.25 35.99
Vanilla Attention Grounding (45k) 95.68 75.45 86.87 53.02 12.42 37.51
Simplified Attention Grounding: GUI-AIMA-3B (45k)
[…]
+ weighting uniformly 97.08 78.34 88.92 56.50 13.91 40.23
+ weighting with all $\mathcal{Q}$ Eq. 5 96.24 78.34 88.44 52.61 13.41 37.63
+ weighting with <ANCHOR> in Eq. 6 96.66 76.35 87.81 56.91 14.57 40.73
[…]
+ layer-wise $\operatorname{top-1}\mathcal{Q}_{s}^{l}$ 96.52 81.23 89.86 57.32 15.73 41.43
+ layer-wise $\operatorname{top-3}\mathcal{Q}_{s}^{l}$ 96.66 79.06 88.99 57.83 15.07 41.49
[…]
+ global $\operatorname{top-3}\mathcal{Q}_{s}$ 95.82 70.94 84.98 57.01 14.40 40.73
[…]
     + weighted patch-wise labeling 96.94 80.14 89.62 59.37 15.89 42.76
GUI grounding without Visual-sink Query Token $\mathcal{Q}_{s}$.
For attention-based GUI grounding variants, weighting the query-visual correlation of each head via the cumulative sum of all query’s predictions in Eq. 5 or only the prediction from <ANCHOR> in Eq. 6 leads to more biased attention predictions than uniform weighting, with worse ScreenSpot-v2 performance of both variants and no improvement on ScreenSpot-Pro as shown in “Attention Head Weighting without $\mathcal{Q}_{s}$” part of Section 4.2. This bias comes from the misleading weighting from visual-unrelated tokens for the variants using all $\mathcal{Q}$ and from using the premature <ANCHOR> token for grounding in the later variant. For attention head weighting softly as in Eq. 11, it achieves similar overall performance on both datasets but performs worse on text elements.
The effect of Visual-sink Query Token $\mathcal{Q}_{s}$.
For ablation of using $\mathcal{Q}_{s}$ for $w_{l,h}$, we vary the $\operatorname{K}$ value between 1 and 3 in the $\operatorname{topK}$ selection of $\mathcal{Q}_{s}$ tokens and explore whether differing $\mathcal{Q}_{s}^{l}$ for each MLLM’s layer as in Eq. 10 or using the same $\mathcal{Q}_{s}$ for all layers as in Eq. 9. In Section 4.2, we can observe that the layer-wise manner performs slightly better on the icon-based tasks, but globally-selected uniform $\mathcal{Q}_{s}$ leads to more balanced results on both text and icon sets. Among both manners, $\operatorname{top-1}$ selection remains the overall best performance. These ablations verify the “global $\operatorname{top-1}\mathcal{Q}_{s}$” setting of GUI-AIMA, with 1.90% improvement over “weighting uniformly” on ScreenSpot-Pro.
Patch-wise labeling considering overlapping and distance.
Here, we also justify the overlapping- and distance-aware (Tang et al., 2025) patch-wise labeling for fine-tuning GUI-AIMA in Eq. 2. Based on “global $\operatorname{top-1}\mathcal{Q}_{s}$”, down-weighting partial positive and distanced image patch instead of labeling all patch equally leads to 1.26% further enhancement on ScreenSpot-Pro.
Training efficiency.
In Section 4.2 , we show the convergence tendency of three methods: GUI-AIMA, vanilla attention grounding and GUI-Actor (with extra warm-up stage already) initialized from Qwen-2.5-VL-3B. GUI-AIMA converges fastest and achieves the best final results, with steady improvements. Vanilla attention grounding fluctuates as supervision on all tokens from the pretrained vocabulary impairs the general capacity. And GUI-Actor needs longer training period to customize its extra modules for GUI grounding.
Figure 4: Analysis experiment results on ScreenSpot-pro. Relax@k measures how many previously incorrect offset predictions are recovered when the ground-truth bounding box is expanded by k visual patches along each dimension. Recovered refers to the number of offset predictions corrected by the second step, while Lost refers to the number of predictions that degrade after the second step.
Model Relax@1 Relax@2 Relax@5 Total Recovered Lost Acc.
[…]
Origin 1-step 158 92 56 306 - - 47.06
[…]
Crop-only 123 83 43 249 156 107 50.16
[…]
Figure 5: Model convergence of the 45k ablation dataset on ScreenSpot-Pro benchmark.
4.3 Analysis of Two-step Inference with Zoom-in
In Section 4.2, we present the analysis results of the two-step inference with zoom-in introduced in Section 3.5. The analysis focuses on incorrect offset predictions, where the predicted center points fall outside the ground-truth bounding box by a small degree. Section 4.2 reports the statistics of offset errors for one-step and two-step inference with different offset distance, and the number of recovered predictions as well as the lost originally-correct predictions after applying the second crop-and-zoom-in step. We observe that performing inference on the focused region determined by the first-step prediction already provides more than a 3% improvement (Crop-only), as fewer irrelevant patches act as distracting noise for our patch-wise grounding. And after cropping, zooming in by a factor from 1.5× to 4× significantly reduces offset errors, especially those with slight deviation (Relax@1). It also helps resolving large-offset errors, achieving up to 31% and 32% error reduction for Relax@3 and Relax@5, respectively. From the analysis results, we find that the performance of two-stem inference is not very sensitive with the zoom-in factor, while a 2-time zoom-in provides the best performance.
5 Conclusion
We presented GUI-AIMA, an attention-based, coordinate-free approach to GUI visual grounding that aligns intrinsic multi-head self-attention with patch-wise supervision. With soft, overlap- and center-aware patch labels converted from coordinate-based annotations, GUI-AIMA simplifies the vanilla attention visual grounding via a learnable <ANCHOR> token as the surrogate to aggregate query-to-visual attention heads. To effectively aggregate different attention heads, we first identify visual-sink query tokens based on query-visual similarity between hidden states. Experimental results of GUI-AIMA on ScreenSpot-v2 and the challenging ScreenSpot-Pro benchmarks, showing state-of-the-art performance at the 3B scale and rivals other methods with larger MLLM backbones, using only about 85k training screenshots. Ablations further verify the effectiveness of each design choice, such as anchored attention aggregation, instruction-adaptive head weighting with visual-sink query tokens, and weighted patch labels. In future work, extending GUI-AIMA to more general and complex visual grounding tasks are pending explorations.
Appendix A Analysis for Visual-sink Query Token
Figure 6: Magnitude of normalized visual correlation of query tokens computed from hidden states (“Embedding”) and from multi-head self-attention (“Attention”).
In Fig. 6, we compare normalized distributions of global visual-token correlation $\sum_{l=1}^{L}c_{q_{i}}^{\,l}$ computed from two different manners. “Embedding” denotes the $c_{q_{i}}^{l}$ computed from Eq. 7 based on hidden states. And “Attention” denotes visual-token correlation of $q_{i}$ with $c_{q_{i}}^{l}=\sum_{h\in[H]}\sum_{v_{i}\in\mathcal{V}}\,\mathbf{A}^{l,h}_{q_{i},v_{i}}$ computed from multi-head attentions. From Fig. 6, we can observe different visual-correlation patterns: while the text token with largest magnitude in “Attention” manner sometimes falls into semantic-irrelevant tokens, such as “this”, “embedding” manner shows larger magnitudes on the <|im_end|> token.
The observations above verify the claim in paper that the query-visual pattern discovered in hidden states $\mathbf{H}^{l}$ is not necessarily statistically prevailing among each head’s self-attention matrix, as only a smaller subset of attention heads are “semantic heads” that key on semantic functionality and representation similarity, which is also supported by Elhelo & Geva (2024); Olsson et al. (2022); Voita et al. (2019).
From ablation results in Section 4.2, the global pattern indicated from visual-sink query token computed $\mathcal{Q}_{s}$ from hidden states achieved better performance, supports our selection that complies with the visual-query pattern from hidden states for weighting attention grounding.
Appendix B Implementations of Baselines
Most baselines’ results are taken from the original papers, except GUI-G2-3B which is not reported in the source.
Appendix C Extra details of ANCHOR token Implementations
In Section 3.3, we abbreviate the implementation of $\mathcal{V},\mathcal{Q},\texttt{<ANCHOR\_START>},\texttt{<ANCHOR>},\texttt{<ANCHOR\_END>}]$ as $\mathcal{V},\mathcal{Q},\texttt{<ANCHOR>}]$ for brevity. For the single-area prediction setting in GUI grounding, we explore to expand single <ANCHOR> to multiple <ANCHOR> tokens as $\mathcal{V},\mathcal{Q},\texttt{<ANCHOR\_0>},\texttt{<ANCHOR\_1>},\ldots,\texttt{<ANCHOR\_N>}]$ (abbreviate start and end tokens). However, it turns out to bring no performance gains and merely adds redundancy for the single-region grounding tasks. We will explore the multi-region grounding tasks with disentangled <ANCHOR_n> token for each grounding region as future works.
Appendix D GUI Grounding Examples of GUI-AIMA
We provide the visualizations of GUI-AIMA’s multi-head attention grounding results on ScreenSpot-v2 and ScreenSpot-Pro as follows in Fig. 7, Figs. 8 and 9.
(a) Lock the memo.
(b) View the language & region settings.
(c) Scan QR code.
(d) View settings of cocowhip light.
(e) Add a new page.
Figure 7: Visualization examples of GUI-AIMA’s grounding results on ScreenSpot-v2.
(a) Add long title.
(b) Change color theme in PyCharm.
Figure 8: Visualization examples of GUI-AIMA’s grounding results on ScreenSpot-Pro.
(a) Save a file.
(b) Select the only feature.
Figure 9: Visualization examples of GUI-AIMA’s grounding results on ScreenSpot-Pro.